# Convergence of patterned matrices with random walk entries

### v1

@@ 4.1 Patterned random matrices with randomly stopped walks
$S_{\pi,n}^{\alpha}=\mathbb{E}(S_{L(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}})-S_{L(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}-1})).$
@@
$S_{\pi,n^{\alpha}}=\mathbb{E}(S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1})).$
As the inverse stable subordinator is an increasing process (see [24]), for distinct indexes $r_{j}$’s, the rvs $(S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1})),\dots,(S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1}))$ are mutually independent due to the independent increments property of the process $\{S_{i,j,n^{\alpha}}(t)=n^{-\alpha/2}\sum_{k=1}^{N(n^{\alpha}t)}X_{i,j,k}\}_{t\geq 0}$. Thus,
@@
(i) for $L(i,j)=(\max\{i,j\},\min\{i,j\})$, the matrix $n^{-1/2}A^{\alpha}_{n}(t)$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges to $B^{\alpha}(t)$ (for definition see (4.1)) in polynomial algebra $\mathbb{C}\langle B^{\alpha}(t)\rangle$ with state defined as $\tau((B^{\alpha}(t))^{l})=\lim_{n\rightarrow\infty}\tau_{n}((A^{\alpha}_{n}(t))^{l})$ for all $l\geq 1$;
(ii) If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j|| for $1\leq i,j\leq n$ then $n^{-1/2}A^{\alpha}_{n}(t)$ algebraically converges to a variable whose moments determine the probability law same as that of $W(L^{\alpha}(t))$, where $\{W(t)\}_{t\geq 0}$ is an SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
4.2 Symmetric circulant matrix and time-changed SBM
For $\alpha\in(0,1]$, let $\lambda_{1,n}^{\alpha}(t),\dots,\lambda_{n,n}^{\alpha}(t)$ be the eigenvalues of the random matrix $A_{n}^{\alpha}(t)$, as defined in (4.5) with link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$. Let $U_{n}$ be a uniform rv on $\{1,\dots,n\}$. We consider a random process $\{Y_{n}^{\alpha}(t)\}_{t\geq 0}$ defined as
@@ 4.2 Symmetric circulant matrix and time-changed SBM
$\lambda_{r,n}^{\alpha}(t)=n^{-\alpha/2}S_{0,n}^{\alpha}(t)+2n^{-\alpha/2}\sum_{j=1}^{[n/2]}S_{j,n}^{\alpha}(t)\cos\bigg(\frac{2\pi jr}{n}\bigg),\ 0<\alpha\leq 1,\ r=1,2,\dots,n,$
@@
$\lambda_{r,n}^{1}(t)=n^{-1/2}S_{0,n}(t)+2n^{-1/2}\sum_{j=1}^{[n/2]}S_{j,n}(t)\cos\bigg(\frac{2\pi jr}{n}\bigg),\ r=1,2,\dots,n,$
@@
Let $Y^{\alpha}_{n}(t)$ be as defined in (4.11). Then, the process $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\in[0,1]}$ weakly converges to $\{W(L^{\alpha}(t))\}_{t\in[0,1]}$, where $\{W(t)\}_{t\geq 0}$ is the SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
@@
In view of Remark 4.2, following the proof of Theorem 3.4, it is easy to show that $\{n^{-1/2}Y^{1}_{n^{\alpha}}(t)\}_{t\in[0,1]}$ weakly converges to the SBM $\{W(t)\}_{t\in[0,1]}$. So, $\{(n^{-1/2}Y^{1}_{n^{\alpha}}(t),L^{\alpha}(t))\}_{t\in[0,1]}$ jointly weakly converges to $\{(W(t),L^{\alpha}(t))\}_{t\in[0,1]}$. Thus, from (4.12) and continuous mapping theorem, $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\in[0,1]}$ converges to $\{W(L^{\alpha}(t))\}_{t\in[0,1]}$ in an appropriate Skorokhod topology (for details see [23], Section 4.4). This completes the proof. ∎

### v2

@@ 4 Patterned matrices with randomly stopped walks
$S_{\pi,n}^{\alpha}=\mathbb{E}(S_{L^{d}(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}})-S_{L^{d}(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}-1}))\dots(S_{L^{d}(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}})-S_{L^{d}(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}-1})).$
@@
$S_{\pi,n^{\alpha}}=\mathbb{E}(S_{L^{d}(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L^{d}(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1}))\dots(S_{L^{d}(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L^{d}(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1})).$
As the inverse stable subordinator is an increasing process (see [24]), for distinct indexes $r_{j}$’s, the rvs $(S_{L^{d}(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L^{d}(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1})),\dots,(S_{L^{d}(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L^{d}(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1}))$ are mutually independent due to the independent increments property of the process $\{S_{i,j,n^{\alpha}}(t)=n^{-\alpha/2}\sum_{k=1}^{N(n^{\alpha}t)}X_{i,j,k}\}_{t\geq 0}$. Thus,
@@
(i) for $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$, the matrix $n^{-1/2}A^{\alpha}_{n}(t)$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges to $B^{\alpha}(t)$ (for definition see (2.3)) in polynomial algebra $\mathbb{C}\langle B^{\alpha}(t)\rangle$ with state defined as $\tau((B^{\alpha}(t))^{l})=\lim_{n\rightarrow\infty}\tau_{n}((A^{\alpha}_{n}(t))^{l})$ for all $l\geq 1$;
(ii) If $L^{1}(i,j)=n/2-|n/2-|i-j||$ for $1\leq i,j\leq n$ then $n^{-1/2}A^{\alpha}_{n}(t)$ algebraically converges to a variable whose moments determine the probability law same as that of $W(L^{\alpha}(t))$, where $\{W(t)\}_{t\geq 0}$ is a BM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
4.1 Symmetric circulant and weak convergence to a time-changed BM
For $\alpha\in(0,1]$, let $\lambda_{1,n}^{\alpha}(t),\dots,\lambda_{n,n}^{\alpha}(t)$ be the eigenvalues of the random matrix $A_{n}^{\alpha}(t)$, as defined in (4.3) with link function $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$. Let $U_{n}$ be a uniform rv on $\{1,\dots,n\}$. We consider a random process $\{Y_{n}^{\alpha}(t)\}_{t\geq 0}$ defined as
@@ 4.1 Symmetric circulant and weak convergence to a time-changed BM
$\lambda_{r,n}^{\alpha}(t)=n^{-\alpha/2}S_{0,n}^{\alpha}(t)+2n^{-\alpha/2}\sum_{j=1}^{[n/2]}S_{j,n}^{\alpha}(t)\cos\Big(\frac{2\pi jr}{n}\Big),\ 0<\alpha\leq 1,\ r=1,2,\dots,n,$
@@
$\lambda_{r,n}^{1}(t)=n^{-1/2}S_{0,n}(t)+2n^{-1/2}\sum_{j=1}^{[n/2]}S_{j,n}(t)\cos\Big(\frac{2\pi jr}{n}\Big),\ r=1,2,\dots,n,$
@@
Let $Y^{\alpha}_{n}(t)$ be as defined in (4.9). Then, the process $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\geq 0}$ converges weakly to $\{W(L^{\alpha}(t))\}_{t\geq 0}$, where $\{W(t)\}_{t\geq 0}$ is a BM that is independent of the inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
@@
In view of Remark 4.2, following the proof of Theorem 3.4, it is easy to show that $\{n^{-1/2}Y^{1}_{n^{\alpha}}(t)\}_{t\geq 0}$ converges weakly to a BM $\{W(t)\}_{t\geq 0}$. So, $\{(n^{-1/2}Y^{1}_{n^{\alpha}}(t),L^{\alpha}(t))\}_{t\geq 0}$ jointly weakly converges to $\{(W(t),L^{\alpha}(t))\}_{t\geq 0}$. Thus, from (4.10) and continuous mapping theorem, $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\geq 0}$ converges to $\{W(L^{\alpha}(t))\}_{t\geq 0}$ in an appropriate Skorokhod topology (for details see [23], Section 4.4). This completes the proof. ∎
5 Non-symmetric patterned random matrices
So far we have dealt with patterned matrices that have a symmetric link function. Now we consider matrices with link functions that are not necessarily symmetric. Then the matrix polynomials are formed using the matrices along with their transposes.
The joint algebraic convergence (convergence of the *-distributions) of iid copies of these matrices can be discussed under assumptions similar to those we have imposed earlier. Specifically, the joint convergence of all the matrices considered in the next sections below holds. The specific patterned matrices we consider are the general circulant matrix (the non-symmetric version of the symmetric circulant matrix), and the elliptic matrix (where the $(i,j)$th and the $(j,i)$th entries are correlated, a special case being the iid matrix, the non-symmetric version of the Wigner matrix). Results similar to those derived earlier for the Wigner and the symmetric circulant are given. However, we skip the detailed proofs.
5.1 Circulant matrix and complex BM
The circulant matrix is defined as $C_{n}:=((a_{L^{1}(i,j)}))_{1\leq i,j\leq n}$ where
$L^{1}(i,j):=(j-i+n)\ \text{mod}\,n,\ 1\leq i,j\leq n.$ (5.1)
It is a non-symmetric matrix and hence has complex eigenvalues. Theorem 3.2.1 of [13] states that if the entries are independent with mean 0, variance $1$, and with uniformly bounded $(2+\delta)$th moment for some $\delta>0$, then the ESD of $n^{-1/2}C_{n}$ converges in probability to the complex (bivariate) Gaussian distribution. The authors used Gaussian approximation to prove this result. Under Assumption I, this convergence can be shown to be almost sure. We omit the detailed proof of this.
Now let $C_{k,n}:=n^{-1/2}((a_{k,L^{1}(i,j)}))_{1\leq i,j\leq n}$, $1\leq k\leq p\in\mathbb{N}$ be $p$ independent $n\times n$ circulant matrices whose entries satsify Assumption I. Then for $1\leq k\leq p$ their eigenvalues are given by
$\lambda^{c}_{k,r}\coloneqq n^{-1/2}\sum_{j=0}^{n-1}a_{k,j}\cos\Big(\frac{2\pi rj}{n}\Big)+\iota n^{-1/2}\sum_{j=0}^{n-1}a_{k,j}\sin\Big(\frac{2\pi rj}{n}\Big),\ 0\leq r\leq n-1.$ (5.2)
Define the following joint ESD:
$n^{-1}\sum_{r=0}^{n-1}\mathbb{I}\big(\text{Re}(\lambda^{c}_{k,r})\leq x_{k},\text{Im}(\lambda^{c}_{k,r})\leq y_{k}),k=1,\ldots p\big).$ (5.3)
Then, by an easy extension of the Gaussian approximation arguments (Berry-Essen bounds) given in Bose [13], we have the following lemma. We omit its proof. We recall that a standard complex Gaussian random variable is of the form $N=(X+\iota Y)/\sqrt{2}$ where $X$ and $Y$ are iid standard Gaussian variables.
Lemma 5.1.
The ESD (5.3) converges weakly a.s. to (the distribution of) $(N_{1},\ldots,N_{p})$ where $\{N_{i}\}$ are iid complex (considered as elements of $\mathbb{R}^{2}$) standard Gaussian random variables.
Let us now turn to joint moment convergence. Note that Theorem 9.5.1 of [9] shows that independent scaled symmetric circulant matrices converge to independent Gaussian variables algebraically. We need a similar result for circulant matrices. By adapting the arguments given there and those given earlier in Section 2.4.2, it is not too difficult to show that only pair-matched words survive when we consider the moment of any monomial, but it does not seem easy to identify/characterize the algebraic limit directly. On the other hand, Lemma 5.3 suggests this algebraic limit must also be $(N_{1},\ldots,N_{p})$. Indeed we have the following result.
Lemma 5.2.
Let $\{N_{k}=N_{k,1}+iN_{k,2}\}_{1\leq k\leq p}$, $p\in\mathbb{N}$ be independent complex Gaussian rvs. Let $\{n^{-1/2}C_{k,n}\}_{1\leq k\leq p}$ be $n\times n$ independent circulant matrices with input sequences satisfying Assumption I. Then, as elements of $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace}(\cdot))$, they converge in *-distribution to $\{N_{1},\ldots,N_{p}\}$. That is,
$\lim_{n\rightarrow\infty}\tau_{n}(g(\{n^{-1/2}C_{k,n},n^{-1/2}C_{k,n}^{T}\}_{1\leq k\leq p}))=\mathbb{E}(g(\{N_{k},\bar{N}_{k}\}_{1\leq k\leq p}))$
for every polynomial $g$. Here, $C^{T}$ and $\bar{N}$ denote the transpose and complex conjugate of $C$ and $N$, respectively.
Proof.
We present only an outline of the proof. As circulant matrices commute, any typical monomial in indeterminates $\{C_{k,n}\}_{1\leq k\leq p}$ and $\{C_{k,n}^{T}\}_{1\leq k\leq p}$ is of the following form: $C_{1,n}^{m_{1}}(C_{1,n}^{T})^{l_{1}}\cdots C_{p,n}^{m_{p}}(C_{p,n}^{T})^{l_{p}}$ for some $\{m_{1},\dots,m_{p},l_{1},\dots,l_{p}\}\subset\mathbb{N}\cup\{0\}$. Note that the circulant matrices are simultaneously diagonalizable using the discrete Fourier matrix. Hence,
$n^{-1}\mathbb{E}\mathrm{Trace}\Big((\dfrac{C_{1,n}}{n^{1/2}})^{m_{1}}\big(\dfrac{C_{1,n}^{T}}{n^{1/2}}\big)^{l_{1}}\cdots\big(\dfrac{C_{p,n}}{n^{1/2}}\big)^{m_{p}}\big(\dfrac{C_{p,n}^{T}}{n^{1/2}}\big)^{l_{p}}\Big)=n^{-1}\sum_{r=0}^{n-1}\mathbb{E}\Big(\lambda_{1,r}^{m_{1}}\bar{\lambda}_{1,r}^{l_{1}}\dots\lambda_{p,r}^{m_{p}}\bar{\lambda}_{p,r}^{l_{p}}\Big),$
where $\lambda_{k,0},\dots,\lambda_{k,n-1}$ are eigenvalues of $n^{-1/2}C_{k,n}$ for each $1\leq k\leq p$. By independence of the matrices, this expectation equals
$n^{-1}\sum_{r=0}^{n-1}\mathbb{E}\Big(\lambda_{1,r}^{m_{1}}\bar{\lambda}_{1,r}^{l_{1}}\Big)\cdots\mathbb{E}\Big(\lambda_{p,r}^{m_{p}}\bar{\lambda}_{p,r}^{l_{p}}\Big).$ (5.4)
Consider each factor in the product. By the Gaussian approximation arguments of [13], we know that for every fixed $r$, $(\lambda_{l,r},1\leq l\leq p)$ converges weakly to $(N_{l},\ldots,N_{p})$. Moreover, using Assumption I, any power of these variables is uniformly integrable, and hence their expectations also converge to the corresponding expectations of the limit variables, for every fixed $r$. Indeed, by using an appropriate non-uniform Berry-Esseen bound, this convergence is also uniform over $0\leq r\leq n-1$. That is, for every $\{m_{k},l_{k}\}$,
$\sup_{\begin{subarray}{c}1\leq k\leq p,\\ 0\leq r\leq n-1\end{subarray}}\Big|\mathbb{E}\big(\lambda_{k,r}^{m_{k}}\bar{\lambda}_{k,r}^{l_{k}}\big)-\mathbb{E}\big(N_{k}^{m_{k}}\bar{N}_{k}^{l_{k}}\big)\Big|\to 0\ \ \text{as}\ \ n\to\infty.$
Since (5.4) is an average over $0\leq r\leq n-1$, the limit in (5.4) equals $\prod_{k=1}^{p}\mathbb{E}\big(N_{k}^{m_{k}}\bar{N}_{k}^{l_{k}}\big)$. This proves the result. ∎
Now, with $L^{1}$, $S_{i,j}^{(n)}(t)$, and $S_{i,j,n}^{\alpha}(t)$ as in (5.1), (3.1) and (4.1), respectively, define the circulant matrices with random walk and stopped random walk entries as follows:
$C_{n}(t)$ $\coloneqq((S_{L^{1}(i,j)}^{(n)}(t)))_{1\leq i,j\leq n},$ (5.5)
$C_{n}^{\alpha}(t)$ $\coloneqq((S_{L^{1}(i,j),n}^{\alpha}(t)))_{1\leq i,j\leq n},\ 0<\alpha\leq 1,\ t\geq 0,$ (5.6)
respectively. Then, in view of Lemma 5.2, similar to Theorem 3.3 (ii) and Corollary 4.1 (ii), we have the following result. We need the (standard) complex Brownian motion (CBM). We can write it as $CBM(t)\stackrel{{\scriptstyle d}}{{=}}(B_{1}+\iota B_{2})/\sqrt{2}$, where $B_{1}$ and $B_{2}$ are independent BMs and $\iota:=\sqrt{-1}$. Its time changed variant is defined in the obvious way and will be denoted by $CBM^{\alpha}(\cdot)$. We omit the proof of the next result.
Theorem 5.1.
Under Assumption I, $\{n^{-1}C_{n}(t)\}_{t\geq 0}$ and $\{n^{-(1+\alpha)/2}C_{n}^{\alpha}(t)\}_{t\geq 0}$ converge algebraically (that is, in their *-distribution) respectively to $\{CBM(t)\}_{t\geq 0}$, and $\{CBM^{\alpha}(t)\}_{t\geq 0}$.
Using (5.2), the eigenvalues of $C_{n}(t)$ and $C_{n}^{\alpha}(t)$ are given respectively by
$\beta_{r,n}$ $\coloneqq\sum_{j=0}^{n-1}S_{j,n}(t)\cos\Big(\frac{2\pi rj}{n}\Big)+\iota\sum_{j=0}^{n-1}S_{j,n}(t)\sin\Big(\frac{2\pi rj}{n}\Big),\ 0\leq r\leq n-1,$ (5.7)
$\beta^{\alpha}_{r,n}$ $\coloneqq\sum_{j=0}^{n-1}S_{j,n}^{\alpha}(t)\cos\Big(\frac{2\pi rj}{n}\bigg)+\iota\sum_{j=0}^{n-1}S_{j,n}^{\alpha}(t)\sin\Big(\frac{2\pi rj}{n}\bigg),\ 0\leq r\leq n-1.$ (5.8)
Let $U_{n}$ be a rv independent of everything else and uniformly distributed on $\{0,2,\ldots,n-1\}$. Similar to $Y_{n}(t)$ and $Y_{n}^{\alpha}(t)$ defined in (3.12) and (4.9), define the empirical processes
$Z_{n}(t)$ $:=\beta_{U_{n},n}(t),\ \text{and}\ \ Z_{n}^{\alpha}(t):=\beta^{\alpha}_{U_{n},n}(t).$
Then similar to Theorems 3.4 and 4.3, we have the following result. We omit the proof.
Theorem 5.2.
Under Assumption I, $n^{-1}Z_{n}(\cdot)$ and $n^{-(1+\alpha)/2}Z_{n}^{\alpha}(\cdot)$ converge weakly respectively to $CBM(\cdot)$ and $CBM^{\alpha}(\cdot)$ in suitable Skorokhod topologies.
5.2 Elliptic matrices with CTRW entries
An elliptic variable $e$ with correlation parameter $1\leq\rho\leq 1$ can always be described as (see page 38 of [9])
$e=\sqrt{\dfrac{1+\rho}{2}}s_{1}+\iota\sqrt{\dfrac{1-\rho}{2}}s_{2},$
where $s_{1}$ and $s_{2}$ are two free standard semi-circular variables. A special case is when $\rho=0$ and in that case we get the standard circular variable, $c\stackrel{{\scriptstyle d}}{{=}}(s_{1}+\iota s_{2})/\sqrt{2}$, where $s_{1}$ and $s_{2}$ are two free standard semi-circular variables.
Suppose $B_{1}(t)_{t\geq 0}$ and $B_{2}(t)_{t\geq 0}$ are free FBMs and are free of each other. Let
$E(t):=\sqrt{\dfrac{1+\rho}{2}}B_{1}(t)+\iota\sqrt{\dfrac{1-\rho}{2}}B_{2}(t),\ t\geq 0.$
Then $E(\cdot)$ has free independent increments, $E(t)=0$, and for every $s<t$, $(E(t)-E(s))/\sqrt{t-s}$ is an elliptic variable with parameter $\rho$. The stopped version of $E(\cdot)$ is defined in the obvious way and will be denoted by $E^{\alpha}(\cdot)$.
Suppose the pairs $(X_{i,j},X_{j,i})$ are independent, each component has mean 0, variance $1$, and all variables have uniformly bounded moments, and the pairs $(X_{i,j},X_{j,i})$ have a common correlation $\rho$ for all $i\neq j$. Then the non-symmetric random matrix $((X_{i,j}))_{1\leq i,j\leq n}$ is called an elliptic matrix. The special case $\rho=0$ includes the fully independent matrix whose all elements are independent. The special case $\rho=1$ includes the Wigner matrix, which is symmetric. It is known that independent scaled elliptic matrices (even with different $\rho$’s) converge in *-distribution to free elliptic variables with the corresponding $\rho$’s (see Theorem 11.1.1 of [9]).
Now consider the following assumption which is similar to Assumption I but incorporates dependent variables, though in a limited way.
Assumption II $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ are rvs such that the pairs $(X_{i,j,k},X_{j,i,k})$ are independent, have a common correlation $\rho$, and each component has mean 0, variance $1$, with all moments uniformly bounded.
Using the result cited above, and going through the arguments given earlier for symmetric matrices with random walk entries, we have the following theorem. We omit its proof.
Theorem 5.3.
Consider the elliptic matrices $E_{n}(t):=((S_{i,j}^{(n)}(t)))_{1\leq i,j\leq n},\ t\geq 0$, where $S_{i,j}^{(n)}(t)$ are as defined in (3.1) and $\{X_{i,j,k}\}$ satisfy Assumption II. Then $\{n^{-1}E_{n}(t)\}_{t\geq 0}$ converges in *-distribution to $\{E(t)\}_{t\geq 0}$. The stopped version $n^{-(1+\alpha)/2}E_{n}^{\alpha}(\cdot)$ of $n^{-1}E_{n}(\cdot)$ converges in *-distribution to $E^{\alpha}(\cdot)$. As a consequence, the ESD of any self-adjoint polynomial in $\{n^{-1}E_{n}(t)\}_{t\geq 0}$ or $\{n^{-(1+\alpha)/2}E_{n}^{\alpha}(t)\}_{t\geq 0}$ converges weakly a.s. to the probability distribution of the corresponding polynomial in $E(\cdot)$ or $E^{\alpha}(\cdot)$ respectively.